OpenStudy (anonymous):

and [6.03] What is the value of the x variable in the solution to the following system of equations? 4x + y = 8 x − 2y = 2 0 2 There is no x value as there is no solution to this system x can be any value as there are infinitely many solutions to this system

OpenStudy (anonymous):

Solve the equations. 4x + y = 8 y = 8 - 4x x - 2y = 2 x - 2 = 2y y = (x - 2)/2 y = y, therefore: (x - 2)/2 = 8 - 4x x - 2 = 16 - 8x x + 8x = 16 + 2 9x = 18 x = 18/9 x = 2 That leaves answer B or D. Rule out D by putting x = 0 into the equations: 4(0) + y = 8 y = 8 and 0 - 2y = 2 2y = 2 y = 1 Since y = y, and 8 = 1 is not correct, there are not infinitely many solutions, and since x is not dependent on an exponent, the solution cannot be -2, so the solution is 2 and only 2, so the answer is B.

OpenStudy (anonymous):

wow! thanks so much!

OpenStudy (anonymous):

No problem. Just realized I made a mistake there at the end: 0 - 2y = 2 2y = 2 y = 1 should be: 0 - 2y = 2 -2y = 2 y = -1

OpenStudy (anonymous):

haha